If z=8(cos5π3+isin5π3), then z6 is equal to


1
Expert's answer
2022-06-06T02:44:44-0400
"z^6=8^6(\cos(6(\dfrac{5\pi}{3}))+i\sin(6(\dfrac{5\pi}{3})))"

"=262144"